OpenStudy (anonymous):

Consider these two fractions: 14/20 and 28/44 a) Reduce each fraction to lowest terms. You must show what you divide the numerator and denominator by. b) Are the fractions equivalent? c) Please repeat steps a and b above for 6/15 and 14/35.

OpenStudy (jdoe0001):

start by see if you can divide top and bottom by 2, if not 2 then 3 or 4 or 5 till you can't anymore, then that'd be the "simplified" version

OpenStudy (jdoe0001):

\(\bf \cfrac{14}{20}\quad \textit{divide by 2 both}\implies \cfrac{7}{10}\) 7 is a prime number, thus not divisible, thus that's the simplified version not try \(\bf \cfrac{28}{44}\)

OpenStudy (anonymous):

so did you multiply \[\frac{ 7 }{ 10 }\] by 4

OpenStudy (jdoe0001):

division, you "reduce them", but the divisor you use, has to work for both top and bottom

OpenStudy (jdoe0001):

in \(\bf \cfrac{7}{10}\) I can divide 10 by 2, but can't do that with 7, so no dice I can divide 10 by 5, but can't do that with 7, so no dice can't divide 10 by 3 or 4 or 6 or else... so.... that's as simplified as it gets

OpenStudy (anonymous):

so \[\frac{ 28 }{ 44 }\] would be \[\frac{ 7 }{ 11 }\]